英語 (English)

主題別英語 (Thematic English) (医 ((医 A)1 年))

ポンドクリストファー · 非常勤講師 / 全学共通教育センター 1 単位後期金 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This class gives students the opportunity to develop their English communicative ability through the exploration and discussion of themes relevant to our modern global world. Practice with speaking, reading, writing and listening will be provided through class activities around a textbook and other media.

【授業の概要】 This is a communicative style class. Students will be expected to take part in various pair-work and small group activities such as role-plays, discussions and information-gap tasks. There will be a short homework activity each week that is necessary preparation for the following week’s class.The mid-term and final exams are presentation based.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】In today’s globalized society, English has become the common

language of communication and is necessary to participate in today’s global community. This class requires students to put their skills into practice, and in the process increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

【授業の計画】 1. Introduction to the class, expectations and getting to know each other.

2. Theme 1: General knowledge quiz

3. Introductory presentations

4. Theme 2: The clown doctor

5. Theme 3: The painter and the writer

6. Theme 4: Books and films

7. Theme 5: Planning a trip

8. Mid-term presentations

9. Theme 6: How to behave abroad

10. Theme 7: Hotels with a difference

11. Theme 8: travelling around 12. Theme 9: The world’s favorite food

13. Theme 10: A lifelong passion

14. Theme 11: Writing a biography

15. Final presentations

16. Review

【教科書】New Headway 3rd Edition Intermediate STUDENT BOOK. Liz and

John Soars, Oxford University Press. ISBN-13 9780 1943 87507 【参考書等】 [参考資料] 【成績評価の方法】 Weekly attendance, active participation and homework are an important part of this course. Students are evaluated as follows: Attendance, participation and homework 50%, mid term presentation 15%, final presentation 35%

【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 Reading English books and keeping an English diary are both recommended as great ways to improve your English ability. Above all be afraid to take any opportunity to communicate in English.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221159 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ポンド .

English

Thematic English (医 ((医 A)1 年))

Christopher Pond · Part-time Lecturer / Center for General Education 1 unit 後期金 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This class gives students the opportunity to develop their English communicative ability through the exploration and discussion of themes relevant to our modern global world. Practice with speaking, reading, writing and listening will be provided through class activities around a textbook and other media.

Outline⟩ This is a communicative style class. Students will be expected to take part in various pair-work and small group activities such as role-plays, discussions and information-gap tasks. There will be a short homework activity each week that is necessary preparation for the following week’s class.The mid-term and final exams are presentation based.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ In today’s globalized society, English has become the common language of

communication and is necessary to participate in today’s global community. This class requires students to put their skills into practice, and in the process increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

Schedule⟩

1. Introduction to the class, expectations and getting to know each other.

2. Theme 1: General knowledge quiz

3. Introductory presentations

4. Theme 2: The clown doctor

5. Theme 3: The painter and the writer

6. Theme 4: Books and films

7. Theme 5: Planning a trip

8. Mid-term presentations

9. Theme 6: How to behave abroad

10. Theme 7: Hotels with a difference

11. Theme 8: travelling around 12. Theme 9: The world’s favorite food

13. Theme 10: A lifelong passion

14. Theme 11: Writing a biography

15. Final presentations

16. Review

Textbook⟩ New Headway 3rd Edition Intermediate STUDENT BOOK. Liz and

John Soars, Oxford University Press. ISBN-13 9780 1943 87507 Reference⟩ [参考資料] Evaluation Criteria⟩ Weekly attendance, active participation and homework are an important part of this course. Students are evaluated as follows: Attendance, participation and homework 50%, mid term presentation 15%, final presentation 35%

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ Reading English books and keeping an English diary are both recommended as great ways to improve your English ability. Above all be afraid to take any opportunity to communicate in English.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221159 Contact (Office-Hour, Room, E-mail)⟩

⇒ Pond .